OpenStudy (anonymous):

i need help with this problem?help please:) what will an investment of 20,000 amount to after 34 months if it earns 6.5% p.a compounded monthly?

OpenStudy (anonymous):

Do you know the equation you can plug into

OpenStudy (anonymous):

i think its p=(1+r/4)^4

OpenStudy (tkhunny):

Not bad, but you're missing a few pieces. You have quarterly and you need monthly. You have 4 periods and you need 34. You don't have the initial deposit. \(Accumulation = 20000\cdot\left(1 + \dfrac{0.065}{12}\right)^{34}\)

OpenStudy (anonymous):

do i have to solve for that equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

"Solve"? Not really. It's a calculation problem from there. Maybe "resolve". You need to find all the parts. 20000 the initial payment 4 the numer of periods per year 34 the total number of periods 0.065 the annual interest to be compounded as specified

OpenStudy (anonymous):

so if i solve it and calculate it then i get my final answer